OpenStudy (anonymous):

The Fiedler family has up to $130,000 to invest. They decide that they want to have at least $40,000 invested in stable bonds yielding 5.5% and that no more than $60,000 should be invested in more volatile bonds yielding 11%. How much should they invest in each type of bond to maximize income if the amount in the stable bond should not exceed the amount in the more volatile bond? What is the maximum income?

OpenStudy (anonymous):

Put as much as allowed in the higher yield, so $60,000 at 11% would be best. The other $70,000 could be invested at 5.5%. This will maximize income, left as an exercise for the student.....

OpenStudy (tkhunny):

F = Amount Invested in 5.5% Stability E = Amount Invested in 11% Risky The Fiedler family has up to $130,000 to invest. \(F + E \le 130000\) They decide that they want to have at least $40,000 invested in stable bonds yielding 5.5% \(F \ge 40000\) and that no more than $60,000 should be invested in more volatile bonds yielding 11%. \(E \le 60000\) the amount in the stable bond should not exceed the amount in the more volatile bond \(F \le E\) Income (Annual, I guess.) (F * 0.055) + (E * 0.11)